OpenStudy (anonymous):

If you flip a coin 6 times, what is the probability of getting HH TT HT in that order?

OpenStudy (dumbcow):

each flip is an independent event, to find probabilities of a series of independent events we multiply the probabilities of each event each flip has a 1/2 probability of being heads or tails answer = (1/2)^6 = 1/64